# references.bib

### v2

@@ 3.1 Large extremal number
Suppose that $r\geq 3$, $n\equiv 1\pmod{r}$ is large enough, and $\lfloor n/r\rfloor=(n-1)/r\equiv 0\pmod{k}$ (note there are infinitely many such $n$). Let $G$ be obtained by embedding $U(K_{1,k-1},\lfloor n/r\rfloor)$ in a small part of $T_{n,r}$ and let $H$ be obtained by embedding $U(P_{k},\lceil n/r\rceil)$ in a large part of $T_{n,r}$ and then extending one of the paths by a single edge within the part, so that $e(H)=e(G)+1$. Let $V_{1}\sqcup\cdots\sqcup V_{r}$ and $U_{1}\sqcup\cdots\sqcup U_{r}$ be the corresponding partitions of $V(G)$ and $V(H)$, respectively, so that the trees are embedded in $V_{1}$ and $U_{1}$. Since
$k>\frac{r-1}{2}\left(2+\frac{5}{r-1}-\frac{4}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}-\frac{4}{r-1}\left(\frac{6}{r-1}-\frac{4}{r}\right)}\right)$
we have
$\frac{1}{r-1}k+\frac{4}{r}-\frac{5}{r-1}+\left(\frac{6}{r-1}-\frac{4}{r}\right)k^{-1}>2.$
@@
Consider any copy $K$ of $T_{m(r-1),r-1}$ which appears in $H$ and assume for a contradiction that a forbidden subgraph of the form $F+K$ appears. If $V(K)\subseteq V-U_{1}$, then $N(V(K))=U_{1}$, so no forbidden graph of the form $F+K$ appears, by the definition of $H[U_{1}]$. Otherwise, $V(K)\ni v\in U_{1}$. If $|V(K)\cap U_{1}|\geq 3$, then $N(V(K)\cap U_{1})\cap U_{1}=\emptyset$, hence $V(F)\cap U_{1}=\emptyset$. Then $V(F)$ intersects two different parts $U_{2},U_{3}$, implying that $K$ is contained in the $r-2$ parts $U_{1},U_{4},\ldots,U_{r}$. But one can check $|V(K)\cap U_{i}|\leq m$ since $\Delta(H[U_{i}])\leq 2<m$ so $|V(K)|\leq(r-2)m$, a contradiction. Thus, $|V(K)\cap U_{1}|\leq 2$, so $|V(K)-U_{1}|\geq m(r-1)-2$, and thus $V(K)\cap U_{i}\neq\emptyset$ for each $i$ as $m(r-2)+2<m(r-1)-2$. Hence, $N(V(K))\subseteq U_{1}$, and the previous argument applies. This covers forbidden subgraphs of forms (1) through (5). Finally, if $H$ contains a copy $K$ of $K_{r+2}$, then $|V(K)\cap U_{i}|\leq 1$ for each $i>1$ and $|V(K)\cap U_{1}|\leq 2$. Hence $|V(K)|\leq r+1$, a contradiction. ∎
@@
Consider any copy $K$ of $T_{m(r-1),r-1}$ which appears in $G$. If $V(K)\subseteq V-V_{1}$, then $N(V(K))=V_{1}$, so no forbidden graph of the form $F+K$ appears, by the definition of $G[V_{1}]$. Otherwise, $V(K)\ni v\in V_{1}$. If $|V(K)\cap V_{1}|\geq k$, then $N(V(K)\cap V_{1})\cap V_{1}=\emptyset$, hence $V(F)\cap V_{1}=\emptyset$. Then $V(F)$ intersects two different parts $V_{2},V_{3}$, implying that $K$ is contained in the $r-2$ parts $V_{1},V_{4},\ldots,V_{r}$. But one can check $|V(K)\cap V_{i}|\leq m$ since $\Delta(G[V_{i}])\leq k-1<m$ so $|V(K)|\leq(r-2)m$, a contradiction. Thus, $|V(K)\cap V_{1}|\leq k-1$, so $|V(K)-V_{1}|\geq m(r-1)-k+1$, and thus $V(K)\cap V_{i}\neq\emptyset$ for each $i$ since $m\geq k-1$. Hence $N(V(K))\subseteq V_{1}$, and the previous argument applies. This covers forbidden subgraphs of forms (1) through (5). Finally, if $G$ contains a copy $K$ of $K_{r+2}$, then $|V(K)\cap U_{i}|\leq 1$ for each $i>1$ and $|V(K)\cap V_{1}|\leq 2$. Hence $|V(K)|\leq r+1$, a contradiction. ∎
@@
Let $H^{\prime}\in\mathrm{EX}(n,\mathcal{F})$. Following subsection 2.1 through 15 shows that $H^{\prime}$ is obtained from $K_{W_{1},\ldots,W_{r}}$ by embedding trees in $W_{1}$, where $|w_{i}-w_{j}|\leq 1$ for $i,j>1$ and $w_{1}-2\leq w_{i}\leq w_{1}+1$ for $i>1$. Suppose there are $h$ parts of size $w_{1}-2$ and $r-h-1$ parts of size $w_{1}-1$, so that
@@
Now $n\equiv 1\pmod{r}$, the right-hand side is an integer, and $10\leq h\leq r-$ 1, so we have $h=0$, and in fact $K_{W_{1},\ldots,W_{r}}=T_{n,r}$. By (1) through (5), $H^{\prime}[W_{1}]$ is a disjoint union of trees, each of which is $P_{k+1}$, $P_{k}$, $K_{1,{k-1}}$, or of order smaller than $k$. Moreover, at most one tree is $P_{k+1}$ and such a tree can only exist if there is no $K_{1,k-1}$. From these restrictions and the fact that $\lfloor n/r\rfloor\equiv 0\pmod{k}$ it is clear that $e(H^{\prime}[W_{1}])$ (hence $e(H^{\prime})$) is maximized exactly when $H^{\prime}\simeq H$. ∎
@@ 4 Proof of 1
Since
$\frac{r-1}{2}\left(2+\frac{5}{r-1}-\frac{4}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}-\frac{4}{r-1}\left(\frac{6}{r-1}-\frac{4}{r}\right)}\right)$
is not an integer, it is strictly less than its ceiling $k$. Hence, 1 gives that $c(r)\leq\frac{k-1}{kr}$. Suppose $Q<\frac{k-1}{kr}$. Let $\mathcal{F}$ be any finite family of graphs with $\chi(\mathcal{F})=r+1$ and $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+Qn$ for large enough $n$. Then we have
$k^{\prime}$ $:=k(\mathcal{F})\leq\left\lfloor\frac{1}{1-Qr}\right\rfloor\leq k-1$
@@ 5 Concluding remarks
Theorem 2 gives a condition on the value of $\mathrm{ex}(n,\mathcal{F})$ which guarantees that $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$, when $\chi(\mathcal{F})=r+1\geq 4$. This condition is sharp for many values of $r$. As shown by 2, Theorem 2 does not hold if we allow $r=2$ and $k(\mathcal{F})=3$. We attempted to provide a similar counterexample in the case of $r=2,k(\mathcal{F})=2$ but were unable to do so because $\lambda(H)$ is in fact greater than $\lambda(G)$ by a term of $\Theta(n^{-2})$, where $G$ and $H$ are graphs defined analogously to the ones in subsection 3.2. This fact, combined with the result of [wang2023conjecture], suggested 1. A solution to 1 would be interesting as it potentially requires stronger techniques than those of this paper and those of [wang2023conjecture].
@@
[…]

### v3

@@ 3.1 Large extremal number
Suppose that $r\geq 3$, $n\equiv 1\pmod{r}$ is large enough, and $\lfloor n/r\rfloor=(n-1)/r\equiv 0\pmod{k}$ (note there are infinitely many such $n$). Let $G$ be obtained by embedding $U(K_{1,k-1},\lfloor n/r\rfloor)$ in a small part of $T_{n,r}$ and let $H$ be obtained by embedding $U(P_{k},\lceil n/r\rceil)$ in a large part of $T_{n,r}$ and then extending one of the paths by a single edge within the part, so that $e(H)=e(G)+1$. Let $V_{1}\sqcup\cdots\sqcup V_{r}$ and $U_{1}\sqcup\cdots\sqcup U_{r}$ be the corresponding partitions of $V(G)$ and $V(H)$, respectively, so that the trees are embedded in $V_{1}$ and $U_{1}$. Since $k>\alpha$ we have that
$\frac{1}{r-1}k+\frac{4}{r}-\frac{5}{r-1}+\left(\frac{6}{r-1}-\frac{4}{r}\right)k^{-1}>2$
@@
Consider any copy $K$ of $T_{m(r-1),r-1}$ which appears in $H$ and assume for a contradiction that a forbidden subgraph of the form $F+K$ appears. If $V(K)\subseteq V-U_{1}$, then $N(V(K))=U_{1}$, so no forbidden graph of the form $F+K$ appears, by the definition of $H[U_{1}]$. Otherwise, $V(K)\ni v\in U_{1}$. If $|V(K)\cap U_{1}|\geq 3$, then $N(V(K)\cap U_{1})\cap U_{1}=\emptyset$, hence $V(F)\cap U_{1}=\emptyset$. Then $V(F)$ intersects two different parts $U_{2},U_{3}$, implying that $K$ is contained in the r2r-2 parts $U_{1},U_{4},\ldots,U_{r}$. But one can check $|V(K)\cap U_{i}|\leq m$ since $\Delta(H[U_{i}])\leq 2<m$ so $|V(K)|\leq(r-2)m$, a contradiction. Thus, $|V(K)\cap U_{1}|\leq 2$, so $|V(K)-U_{1}|\geq m(r-1)-2$, and thus $V(K)\cap U_{i}\neq\emptyset$ for each $i$ as m(r2)+2<m(r1)2m(r-2)+2<m(r-1)-2. Hence, $N(V(K))\subseteq U_{1}$, and the previous argument applies. This covers forbidden subgraphs of forms (1) through (5). Finally, if $H$ contains a copy $K$ of $K_{r+2}$, then $|V(K)\cap U_{i}|\leq 1$ for each $i>1$ and $|V(K)\cap U_{1}|\leq 2$. Hence $|V(K)|\leq r+1$, a contradiction. ∎
@@
Consider any copy $K$ of $T_{m(r-1),r-1}$ which appears in $G$. If $V(K)\subseteq V-V_{1}$, then $N(V(K))=V_{1}$, so no forbidden graph of the form $F+K$ appears, by the definition of $G[V_{1}]$. Otherwise, $V(K)\ni v\in V_{1}$. If $|V(K)\cap V_{1}|\geq k$, then $N(V(K)\cap V_{1})\cap V_{1}=\emptyset$, hence $V(F)\cap V_{1}=\emptyset$. Then $V(F)$ intersects two different parts $V_{2},V_{3}$, implying that $K$ is contained in the r2r-2 parts $V_{1},V_{4},\ldots,V_{r}$. But one can check $|V(K)\cap V_{i}|\leq m$ since $\Delta(G[V_{i}])\leq k-1<m$ so $|V(K)|\leq(r-2)m$, a contradiction. Thus, $|V(K)\cap V_{1}|\leq k-1$, so $|V(K)-V_{1}|\geq m(r-1)-k+1$, and thus $V(K)\cap V_{i}\neq\emptyset$ for each $i$ since $m\geq k-1$. Hence $N(V(K))\subseteq V_{1}$, and the previous argument applies. This covers forbidden subgraphs of forms (1) through (5). Finally, if $G$ contains a copy $K$ of $K_{r+2}$, then $|V(K)\cap U_{i}|\leq 1$ for each $i>1$ and $|V(K)\cap V_{1}|\leq 2$. Hence $|V(K)|\leq r+1$, a contradiction. ∎
@@
Let $H^{\prime}\in\mathrm{EX}(n,\mathcal{F})$. Following subsection 2.1 through 15 shows that $H^{\prime}$ is obtained from $K_{W_{1},\ldots,W_{r}}$ by embedding trees in $W_{1}$, where $|w_{i}-w_{j}|\leq 1$ for $i,j>1$ and $w_{1}-2\leq w_{i}\leq w_{1}+1$ for $i>1$. Suppose there are $h$ parts of size $w_{1}-2$ and rh1r-h-1 parts of size $w_{1}-1$, so that
@@
Now $n\equiv 1\pmod{r}$, the right-hand side is an integer, and $0\leq h\leq r-1$, so we have $h=0$, and in fact $K_{W_{1},\ldots,W_{r}}=T_{n,r}$. By (1) through (5), $H^{\prime}[W_{1}]$ is a disjoint union of trees, each of which is $P_{k+1}$, $P_{k}$, $K_{1,{k-1}}$, or of order smaller than $k$. Moreover, at most one tree is $P_{k+1}$ and such a tree can only exist if there is no $K_{1,k-1}$. From these restrictions and the fact that $\lfloor n/r\rfloor\equiv 0\pmod{k}$ it is clear that $e(H^{\prime}[W_{1}])$ (hence $e(H^{\prime})$) is maximized exactly when $H^{\prime}\simeq H$. ∎
@@ 4 Proof of 1
First we show that $\alpha$ is not an integer. We have
$\alpha$ $=\frac{r-1}{2}\left(2+\frac{5}{r-1}-\frac{4}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}-\frac{4}{r-1}\left(\frac{6}{r-1}-\frac{4}{r}\right)}\right)$
$=r-1+\frac{5}{2}-2+\frac{2}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}\frac{(r-1)^{2}}{4}-(r-1)\left(\frac{6}{r-1}-\frac{4}{r}\right)}$
$=\frac{2r(r-1)+5r-4r+4}{2r}+\sqrt{\left(\frac{2r^{2}-r+2}{2r}\right)^{2}-\frac{5}{r}+\frac{3}{r^{2}}}.$
For $i\in\mathbb{Z}$ let $s_{i}=(i/2r)^{2}$ and let $S=\{s_{i}:i\in\mathbb{Z}\}$. Note that $s_{2r^{2}-r+2}-s_{2r^{2}-r+1}=1-\frac{1}{2r}-\frac{3}{4r^{2}}.$ Whenever $r\geq 6$, we have
$0<\frac{5}{r}+\frac{3}{r^{2}}<1-\frac{1}{2r}-\frac{3}{4r^{2}}$
so the expression in the square root above is not an element of $S$ and therefore $\alpha$ is not an integer. When $r=3,4,5$, one may check directly that $\alpha$ is not an integer.
Now let $k=\lceil\alpha\rceil>\alpha$. 1 gives that $c(r)\leq\frac{k-1}{kr}$. Suppose $Q<\frac{k-1}{kr}$. Let $\mathcal{F}$ be any finite family of graphs with $\chi(\mathcal{F})=r+1$ and $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+Qn$ for large enough $n$. Then we have
$k^{\prime}$ $=k(\mathcal{F})\leq\left\lfloor\frac{1}{1-Qr}\right\rfloor\leq k-1$
@@ 5 Concluding remarks
Theorem 2 gives a sharp condition on the value of $\mathrm{ex}(n,\mathcal{F})$ which guarantees that $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$, when $\chi(\mathcal{F})=r+1\geq 4$. As shown by 2, Theorem 2 does not hold if we allow $r=2$ and $k(\mathcal{F})=3$. We attempted to provide a similar counterexample in the case of $r=2,k(\mathcal{F})=2$ but were unable to do so because $\lambda(H)$ is in fact greater than $\lambda(G)$ by a term of $\Theta(n^{-2})$, where $G$ and $H$ are graphs defined analogously to the ones in subsection 3.2. This fact, combined with the result of [wang2023conjecture], suggested 1. A solution to 1 would be interesting as it potentially requires stronger techniques than those of this paper and those of [wang2023conjecture].
@@
6 Acknowledgements
The author would like to thank Michael Tait for suggesting the problem and for useful discussions which improved the quality of the paper. This research was partially supported by NSF DSM-2245556.